0/6.66 points I AE.094 Ask Your Teacha binary compound created by reaction of an unknown element E and oxygen contains 30.09% E and 61.11% 0 by mass. ห the formula ofthe compound ist,o,. calculate the atomic mass of E. -27 x amu
Answer #1

Mass E = 38.89 g
Mass O = 61 .11 g => 61.11 / 16 g/mol =3.82 mol
2 : 3= x : 3.82
x = 2.546 moles E
atomic mass E = 38.89 / 2.546=15.27 g/mol

Do give positive rating if this answer is heldhe you